Gucchi:

Triangle JKL is reflected over the y-axis and then dilated by a scale factor of 2 centered at the origin. Which of the following statements describes the properties of triangles JKL and J''K''L'' after the transformations? Corresponding angles are congruent after the reflection, but not after the dilation. Corresponding angles are congruent after both the dilation and the reflection. Corresponding sides are congruent after the dilation, but not after the reflection. Corresponding sides are congruent after both the reflection and the dilation.

Gucchi:

@jhonyy9

Gucchi:

am i right with A?

jhonyy9:

reflection and dilation do change angle measure ?

jhonyy9:

no never

Gucchi:

@jhonyy9 wrote:
no never
yes right

Gucchi:

they dont change

jhonyy9:

so in this way what choice will be right ?

Gucchi:

so its either B or D

Gucchi:

But i would say D because reflection comes first right?

jhonyy9:

when you dilate a side get the same side length ?

Gucchi:

@jhonyy9 wrote:
when you dilate a side get the same side length ?
Yes

jhonyy9:

what mean dilatation ? dilate a side of 2 by 2 so result a side of 2×2= 4 so now what you think about this ?

Gucchi:

WAIT B is correct.

Gucchi:

Corresponding angles are congruent after both the dilation and the reflection.

jhonyy9:

please thought about the right answer

jhonyy9:

think on it

Gucchi:

@jhonyy9 wrote:
what mean dilatation ? dilate a side of 2 by 2 so result a side of 2×2= 4 so now what you think about this ?
so you would get a different side length

jhonyy9:

@gucchi wrote:
@jhonyy9 wrote:
what mean dilatation ? dilate a side of 2 by 2 so result a side of 2×2= 4 so now what you think about this ?
so you would get a different side length
perfect

Gucchi:

So A would be correct since dilation would give you a different side length?

Gucchi:

@jhonyy9 are you here?

jhonyy9:

the second part of A do you think it true ?

Gucchi:

no it doesnt look true

Gucchi:

The answer is definitely B

jhonyy9:

@gucchi wrote:
The answer is definitely B
perfect

Gucchi:

thank you so much

jhonyy9:

anytime